OpenStudy (andrejrusa12):

Which of the following is a solution to the equation 2/3 x + 16 = -14?

OpenStudy (andrejrusa12):

a. -45 b. -30 c. 20 d. -3

OpenStudy (andrejrusa12):

Both if you can :)

OpenStudy (partyshotz2409):

ok

OpenStudy (andrejrusa12):

Thxs lol, for not leaving :)

OpenStudy (partyshotz2409):

So first we have to get X by itself

OpenStudy (andrejrusa12):

ok

OpenStudy (evoker):

whoa there if you are going to do the 2/3 inverse part, you need to do it to all three parts much easier to the 17 actually.

OpenStudy (evoker):

*-16 part first.

OpenStudy (evoker):

fine then you need to multiply all three parts by 3/2

OpenStudy (evoker):

so x+ 16*3/2=-14*3/2

OpenStudy (evoker):

or x+24=-21

OpenStudy (andrejrusa12):

yea but thats not one of the options?

OpenStudy (evoker):

well finish the problem, so x= -21-24 or x=-45

OpenStudy (andrejrusa12):

thxs<3

OpenStudy (evoker):

much easier by the way to do the subtraction first and then undo the fraction bit in this case.